OpenStudy (fools101):

In 2008, Joe bought a new truck for $34 thousand. It depreciates 13% each year If this model holds true, what was the truck worth in 2009? A) $29.6 thousand B) $25 thousand C) $41.2 thousand D) none of the above In 2008, Joe bought a new truck for $34 thousand. It depreciates 13% each year If this model holds true, what was the truck worth in 2009? A) $29.6 thousand B) $25 thousand C) $41.2 thousand D) none of the above @Mathematics

OpenStudy (fools101):

Just say yes or no I will do it my self I don’t want to be annoying just try to finish math : / Thank u for ur help!!!

OpenStudy (fools101):

i think it C?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

try to solve it !

OpenStudy (anonymous):

If the truck depreciates it will be worth less in 2009 than it was in 2008. Depreciates means the value is decreasing...

OpenStudy (anonymous):

let us know what you get. Show your work

OpenStudy (fools101):

oh what this make no sense to me sorry i'm really am struggling ....

OpenStudy (anonymous):

Hmmmmm. Your issue seems to be some confusing with computing the percents of quantities...

OpenStudy (anonymous):

If I ask you what is 60% of $80, do you know how to solve this?

OpenStudy (fools101):

so it would be $20.. right..

OpenStudy (anonymous):

First, convert the percent to a decimal. To do this, just move the decimal place over to the LEFT two places. 60% becomes 0.60.

OpenStudy (anonymous):

Then to find 60% of $80...mutiply the decimal form of the percent (0.60), by $80. So 60% of $80 is:\[0.60*80=$48\]

OpenStudy (anonymous):

Now, try finding 55% of $40.

OpenStudy (fools101):

so it would be 55%X$40

OpenStudy (anonymous):

First, notice that the answer better be close to $20. Why? Well, $20 is half (or, 50%) of $40. 55% is not much different than 50%...

OpenStudy (anonymous):

Remember you have to convert the percent to a decimal first

OpenStudy (fools101):

and it come to be 22.00 right..

OpenStudy (fools101):

like that..

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

Ok...now that you know how to find percents of quantities you can do this problem.

OpenStudy (anonymous):

We are starting with $34 (thousand). It depreciates (decreases) by 13% every year.

OpenStudy (anonymous):

One year has passes in going from 2008 to 2009, so the value depreciates by 13% in this time.

OpenStudy (anonymous):

If I ask you how much it has depreciated by I am asking you "how much is 13% of $34 (thousand)"?

OpenStudy (fools101):

it would be 44.2?!

OpenStudy (anonymous):

Well, it can't be 44.2 because that would mean the truck has gained in value from $34. It is depreciating so our answer must be less than $34

OpenStudy (anonymous):

What is 13% of $34?

OpenStudy (anonymous):

Remember to move the decimal place TWO spots to the left to get your decimal form

OpenStudy (fools101):

i'm still X?! right..

OpenStudy (fools101):

so it would be 4.42

OpenStudy (anonymous):

Yeah!

OpenStudy (anonymous):

So our trucks value has decrease by $4.42. It was $34...what is it now?

OpenStudy (anonymous):

Another way to say this is: you had $34 and I took away $4.42. How much do you have left?

OpenStudy (fools101):

then it would be $29.58

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

there is an easier way to do this. Starting with $34 and taking away 13%, you are left with 1-0.13=0.87 (87%) of what you started with. So you can solve this in one step by computing 87% of $34. 0.87*34=29.58

OpenStudy (fools101):

AWWW so it would be A!? right!

OpenStudy (anonymous):

Yep, A it is

OpenStudy (fools101):

OH SWEET thank u for taking time to help me, it helps a lot when people help other people it means so much to me thank u!!!!!

OpenStudy (anonymous):

The answer is B

OpenStudy (anonymous):

Yeah, I'm not agreeing with answer b

OpenStudy (fools101):

While thank u! : )!

OpenStudy (anonymous):

np:)